# Unveiling Molecular Moieties through Hierarchical Graph Explainability

### 29 Jan 2024 v1

@@ 4 Discussion
In the light of the presented results, The method could be considered as a guidance, together with classical structure-based methods, to analyze crucial molecular moieties involved in the successful binding to a specific target. As showed in fig. 5, our methodology outperform the classical fingerprints-input approach for target assignement. It is well known, in fact, that molecular fingerprints can suffer for the correct interpretation of molecular moieties, especially for large and diverse chemical spaces datasets [25]. The combined use of different levels of molecular embeddings seems to be more and more efficient in catching crucial molecular features for the target assignment. The main strength of the presented approach is the combination of different layers of molecular embeddings (from single atom to complex structures). Every layer is capable to catch different aspects of a molecule, and every step enriches the others giving a complete interpretation of the molecule and of the moieties useful to interact with a specific target. As demonstrated in Fig. 4 for Dinaciclib, the complementarity of the layers is crucial in order not to losse chemical information. When going deeper on the explainability part of our method, it is clear that the proposed model is capable to rationally explain ligand moieties involved in the stabilisation of the host-guest interaction. For the majority of ligands analysed (17 out of 19), our method was in accordance with what observed using protein ligand complexes from experimental crystals or docking algorithms (when crystals were not available). In details, when the message passing mechanisms arrives to the third level, the entire structure is considered, accounting for ’regions’ that actively contribute to the classification. Considering the whole structure of the molecules, instead of single atoms or small patterns really gives ideas about the important roles of the molecules to reach the target and bind to it strongly. For two compounds analysed, the method did not found relevant chemical parts to bind the CDK1 receptor. Given the nature of DL techniques used in this work, it is not possible to conclude why for these molecules the model failed. But for sure chemical structures of Flavopiridol and RGB286638 differ from the other structures. For the majority of molecular data in the training set, H-bond are mainly driven by nitrogen instead of oxygen, so probably the flavonoid structure of the FLavopiridol, is not easily read and interpreted by our network. Indeed, the molecular complexity of RGB286638 probably let the molecule to fall out of the applicability domain of the training and test set. For that reason, our model could fail in retrieving functional parts as interacting core. Similarly, Flavopiridol, a flavonoid derivatives, has a non common structure when compared to those in the initial dataset, while manatining the h-bond acceptor/donor pattern useful for CDK1 binding.
@@ 5 Conclusion
In the last years, Artificial Intelligence has been largely adopted in drug design virtual screening campaigns in order to strenghten predictions. The possibility of training models on known actives and inactives has been demonstrated to be a valuable tool to reduce false positives and catch molecular properties fundamental for ligand-protein interactions. The use of AI in chemoinformatics allowed the rise of ligad-based methods that were outclassed by structure-based methods in the last years. Structure-based methods offer the possibility to look deeper in the binding site of a protein allowing scientists to predict how ligand could bind to a specific target. Anyway, It is well known that the use of structure-based methods in drug design could be strongly affected by the protein and ligand conformation in retrieving important protein-ligand interactions. Moreover, high quality protein-ligand complexes are not always available in public databases. Differently to structure-based methods, the results obtained from our method, enlighten not only the interacting partners but also the molecular moieties involved, for example, in the hydrophobic interactions (useful for entropy driven binding) and other chemical parts responsible for the recognition of CDK1. The main strenght of our method is the use of combined molecular embeddings layers that taken together are capable to account the whole molecular structure and functional groups connections. As a result of this architecture, it is possible to correctly assign a target to each molecule and retrieve the main molecule moieties involved in the target assignation/binding. Such a data is not retrievable in other computational techniques used for virtual screening, letting our method to unveil hidden molecular moieties useful for binding a specific target. Our methods herein presented, will be applied to other targets and molecules prioritised by graph-ember will be tested on a panel of kinases in order to experimentally validate the approach.
@@
A detailed description of the predictive results obtained from the Hierarchical Gradient Explainer, (HGE) approach for all 19 approved CDK1 drugs is provided in the attached file “supplementary_material.pdf”.

### 29 Feb 2024 v2

@@ 4 Discussion
In view of the presented results, our explainer could be considered as a guidance, together with classical structure-based methods, to analyze crucial molecular moieties involved in the successful binding to a specific target. We showed that our multi-target GC classifier outperforms the EMBER classifier, proposed by some of the authors, for target assignment. It is well known, in fact, that molecular fingerprints can suffer for the correct interpretation of molecular moieties, especially for large and diverse chemical spaces datasets [25]. The combined use of different levels of molecular embeddings seems to be more and more efficient in catching crucial molecular features for assigning the target. The main strength of the presented approach is the combination of different explanation layers (from single atoms to complex structures). Every layer is capable to catch different aspects of a molecule, and every step enriches the others giving a complete interpretation of both the molecule and the moieties useful to interact with a specific target. As it is demonstrated in Figure 6 for Dinaciclib, all the layers complement each other in a crucial way to avoid loosing chemical information. When going into more detail, it is clear that the proposed model is capable to rationally explain ligand moieties involved in the stabilisation of the host-guest interaction. For the majority of the investigated ligands (17 out of 19), our method was in accordance with the observations resulting from the use of either protein-ligand complexes from experimental crystals or docking algorithms (when crystals were not available). When the message passing mechanism at the topmost level is regarded, the entire structure is considered, accounting for ‘regions’ that actively contribute to the classification. Considering the whole structure of the molecules instead of single atoms or small patterns, provides insights about the important roles of the molecules to reach the target and bind to it strongly. For two compounds analysed, the proposed explainer did not found relevant chemical parts to bind the CDK1 receptor. Given the nature of the DL techniques used in this work, it is not possible to devise exactly the reason of the model’s failure for these molecules. For sure, the chemical structures of Flavopiridol and RGB286638 differ from the other molecules in the training set where the H-bonds are mainly driven by nitrogen instead of oxygen. Indeed, one can think that the molecular complexity of RGB286638 puts this ligand out of the applicability domain of the training and test set. For this reason, our model could fail in detecting some functional parts of the molecule as the interacting core. Similarly, Flavopiridol, a flavonoids derivative, has a non common structure when compared to those in the initial dataset, while maintaining the H-bond acceptor/donor pattern useful for CDK1 binding.
@@ 5 Conclusion
The use of AI in chemoinformatics allowed the rise of ligand-based methods that were outclassed by structure-based methods in the last years. Structure-based methods offer the possibility to look deeper in the binding site of a protein allowing scientists to predict how ligand could bind to a specific target. Anyway, It is well known that the use of structure-based methods in drug design could be strongly affected by the protein and ligand conformation in retrieving important protein-ligand interactions. Moreover, high quality protein-ligand complexes are not always available in public databases. In our work, a novel mixed ligand/structure-based VS approach has been presented in this work, which exploits structural information using a novel eXplainable AI (XAI) procedure on a Deep Neural Network trained for a ligand-based VS task. The presented explainer leverages the algorithmic characteristics of the well-known Grad-CAM explainer for Graph Convolutional Neural Networks based on message passing. Moving from the SOTA results obtained in ligand-based VS with a multi-target classifier which uses the EMBER molecular descriptor, we built a novel classifier relying on GCNNs applied to molecular graphs. Such a multi-target classifier outperformed our previous result, and formed the trained backbone for our explainability analysis. The main idea is using a hierarchy of Grad-CAM explainers placed at different layers of the network where the message passing takes into account a wider and wider neighborhood of each atom. The explanations at each level are then summed together to provide a global map where the different molecular moieties involved in the bioactivity prediction are highlighted and labelled with their role using the well known RDKit package. In our implementation, we applied the explainer to a single target version of the classifier to avoid the inherent classification bias arising in multi-class architectures. We built a CDK1-only classifier that was tuned properly, and the explainability results on 19 well known drugs with CDK1 inhibitory action were tested by a Computational Chemistry expert. The results are very good: 17 of 19 explanations are in accordance with the docking analysis carried out by the expert.
Differently from pure structure-based methods, the results obtained from our method enlighten not only the interacting partners but also the molecular moieties involved, for example, in the hydrophobic interactions (useful for entropy driven binding) and other chemical parts responsible for the recognition of the target. The main strength of our method is the use of combined explainability layers that taken together are capable to account for both the whole molecular structure and functional groups connections. As a result of this architecture, it is possible to correctly assign a target to each molecule and retrieve the main molecule moieties involved in the target assignation/binding. Such a result is not retrievable in other computational techniques used for Virtual Screening, letting our method to unveil hidden molecular moieties useful for binding a specific target. Our methods The explainability approach herein presented, will be applied to other targets, and the molecules prioritised by the multi-target classifier will be tested on a panel of kinases in order to experimentally validate the approach.
@@
A detailed description of the predictive results obtained from our hierarchical explainer for 19 approved CDK1 drugs is provided in the attached file “supplementary_material.pdf”.